A class consisting of 30 girls and 40 boys sponsored a Halloween dance. If 60 percent of the girls and 25 percent of the boys went to the dance, what percent of the entire class went to the dance?
a)30 % b) 40 % c) 45 % d) 70 % e) 85 %

Answer by richwmiller(17219)

3/5*30=18
1/4*40=10
10+18=28/70=4/10=40%